Use the Substitution method to solve the system of equations.
y - 2x = -5
3y - x = 5

Answer by checkley77(12844)

3x + y = 5
y=-3x+5
now replace y with(-3x+5) for x in the other equation & solve for x.
4x - 7y = -10
4x-7(-3x+5)=-10
4x+21x-35=-10
25x=-10+35
25x=25
x=25/25
x=1 answer.
y=-3*1+5
y=-3+5
y=2 answer.
proof
4*1-7*2=-10
4-14=-10
-10=-10